Circular No. 5783

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993M IN ANONYMOUS GALAXY J. Maza, University of Chile; and M. Hamuy, Cerro Tololo Interamer-

ican Observatory, report the discovery of an apparent supernova of m_{pg} ∼ 18.5 by R. Antezana (University of Chile) on a 20-min unfiltered IIa-O plate taken by G. Valladares with the CTIO Curtis Schmidt telescope on Apr. 28.33 UT. The new object is located 1′′ west and 3′′ south of the nucleus of a galaxy at α = 19^{h}08^{m}16^{s}, δ = −64^{o}22′29′′ (equinox 1950.0). Confirmation of SN 1993M was made by R. Aviles on CCD B, V, and I images obtained on May 2.279 with the CTIO 0.9-m telescope.

SUPERNOVA 1993J IN NGC 3031 J. Y. Hu, L. F. Wang, and X. J. Jiang, Beijing Astronomical Observa-

tory, report: “Further observations and analyses of the spectroscopic data of SN 1993J obtained at the Beijing Astronomical Observatory’s 2.16-m telescope show that the profile of Hβ emission line is also double-peaked, in a way similar to that reported on IAUC 5779 for Hα. The local minimum is at about the rest wavelength in our low-resolution spectra (0.53 nm/pixel). The two peaks are shifted by ± 3600 km/s for both the Hα and Hβ profiles. These features imply an asymmetric geometry (cf. IAUC 5777, 5780). Preliminary modelling shows that axially-symmetric structures such as a torus can fit the line profiles. Two unidentified emission features at 695.0 and 704.0 nm appeared on Apr. 22 and are getting more prominent in the spectra taken afterwards. It is not clear if these are ‘satellite’ features of the Hα line.”

Further photoelectric photometry (cf. IAUC 5750, 5755, 5758) by D. Hanzl, N. Copernicus Planetarium, Brno (H); by M. Kato, Ouda Station, Kyoto University, Nara (MK); by V. Karlovsky, Hlohovec, Slovak Republic (VK; comparison star 24 UMa); and by P. Prugniel, Observatoire de Haute Provence (P): Apr. 20.91 UT, V = 10.89 ± 0.02, B–V = +0.68 ± 0.03 (VK); 25.62, V = 11.27 (MK); 25.64, I_{c} = 10.60 (MK); 25.89, V = 11.34 ± 0.07, B–V = +1.08 ± 0.07 (H); 26.92, V = 11.39 ± 0.03, B–V = +1.13 ± 0.05 (H); 29.00, V_{c} = 11.62, B_{c}–V_{c} = +1.25, V_{c}–R_{c} = +0.56, V_{c}–I_{c} = +0.79 (P); May 1.07, V_{c} = 11.74, B_{c}–V_{c} = +1.27, V_{c}–R_{c} = +0.59, V_{c}–I_{c} = +0.84 (P); 3.00, V_{c} = 11.86, B_{c}–V_{c} = +1.30, V_{c}–R_{c} = +0.60, V_{c}–I_{c} = +0.88 (P); 4.00, V_{c} = 11.90, B_{c}–V_{c} = +1.29, V_{c}–R_{c} = +0.60, V_{c}–I_{c} = +0.88 (P).

1993 May 4 Daniel W. E. Green